HUBBLE SPACE TELESCOPE DAILY REPORT #1058 PERIOD COVERED: 0700 EST Daily Status Report as of 021/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FOS/BL 4745 (COSTAR/FOS Fine Alignment - Focus) The COSTAR M1 mirror was successfully moved to a new location. The subsequent observation showed no problem except that the final peakup for the Red Side observations failed to execute. The FOS shutter opened, but the dwell scan never occurred. There was a loss of lock which began at the same time that the shutter opened. An HSTAR was written. Subsequent observations were executed as scheduled. A long series of observations were executed in each orbit with a small angle maneuver, a peakup acquisition, and two images. One of the observations showed low throughput, and the others were reported to be nominal. 1.2 Completed S/C 5254 (GHRS SMOV Side 1 Activation) The Side One DEB was successfully turned on, at three hours intervals. The GHRS low voltage came on Side One without apparent difficulty. 1.3 Completed WF/PC-2 4763 (WF/PC-2 QE and Throughput Measurement) The WF/PC-2 was used to make ultraviolet exposures of the UV flux-standard star BD+75D325 to be used for alignment studies. The OSS monitored the observations, and assessed the data quality. A total of 11 images were received and displayed in the OSS. No anomalies were noted. 1.4 Completed Two Sets of WF/PC-2 5240 (SMOV-WF/PC-2 Flat Fields Calibration) The WF/PC-2 was used to make a set of visual internal flat images. 11 images were received and displayed. Three images were reported to be underexposed, and the others were okay. 1.5 Completed FOC/96 4788 (SMOV COSTAR/FOC Point Spread Function) Both the Faint Object Camera and the COSTAR were used to measure the system's point spread function with high photometric accuracy using ultraviolet standard star HZ4. A corrective maneuver was successfully calculated and uplinked to the spacecraft. Subsequent observations were executed as scheduled. No anomalies were noted. 1.6 Completed COSTAR/U 4814 (COSTAR GHRS M2 Mirror Arm Deployment) The GHRS M2 arm on the COSTAR was successfully deployed as scheduled. The "first light" for the High Resolution spectrograph is expected to happen tonight. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 6 Scheduled Reacquisitions : 16 Successful : 16 2.2 FHST Updates: Scheduled :9 Successful :9 2.3 Significant HSTARs: None 2.4 Operations Notes: All HST subsystems are performing nominally, as SMOV activities continue to be done in accordance with the planned schedule. The GHRS M2 arm was deployed. This instrument also successfully executed tilt and focus operations in preparation for "first light" tomorrow. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete FGS 4892 (Duplicity Among Hyades Stars, Cycle 3 High) 3.2 Complete Seven Sets of WF/PC-2 5240 (SMOV-WF/PC-2 Flat Field Calibration) 3.3 Complete COSTAR/U 4745 (COSTAR/FOS Fine Alignment - Focus) 3.4 Complete S/C 4616 (WF/PC-2 Fine Alignment and focus) 3.5 Complete Three Sets of S/C 5254 (GHRS SMOV Side One Activation) 3.6 Complete Two Sets of FOS/RD 4908 (SMOV FOS/FGS Fine Alignment) 3.7 Complete Two Sets of WF/PC-2 4819 (WF/PC-2 PSF Characterization) 3.8 Complete GHRS/2 4891 (GHRS/FGS Coarse Alignment) 3.9 Complete WF/PC-2 4721 (WF/PC-2 Decontamination and TEC Turn On) 3.10 Complete FGS 5057 (Extragalactic Astrometry and Astrophysics, FT0-11, Part 1 of 2) 3.11 Complete FGS 5063 (Linking the HIPPARCOS to the Extragalactic Reference Frame, Part 5 of 6) 3.12 Complete Two Sets of WF/PC-2 4749 (WF/PC-2 Relative Photometric Calibration) 3.13 Complete WF/PC-2 4709 (WF/PC-2 Internal Monitor) 3.14 Complete FOC/.96 4700 (FOC Flat Field Check-out) 3.15 Complete FOC/96 4788 (SMOV COSTAR/FOC Point Spread Function) 3.16 Complete S/C 5582 (FOC/48 Monthly Activation (camera Section Only) Cycle 4) 3.17 Complete WF/PC-2 4763 (WF/PC-2 QE and Throughput Measurement) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: None /DMH -END- C